Find the 30th term of the series: 15, 18.5, 22…. and the sum of the first 30 terms of the same series.
Answer #1

This is an arithmetic progression

First term = 15, common difference = 3.5

30th term : 15 + (30 - 1) x 3.5 = 116.5

Sum of first 30 terms = 30/2 x (15 + 116.5) = 1972.5